# CMHSU: An R Statistical Software Package to Detect Mental Health Status, Substance Use Status, and their Concurrent Status in the North American Healthcare Administrative Databases

### v3

@@ Abstract
The concept of concurrent mental health and substance use (MHSU) and its detection in patients has garnered growing interest among psychiatrists and healthcare policymakers over the past four decades. Researchers have proposed various diagnostic methods, including the Data-Driven Diagnostic Method (DDDM), for the identification of MHSU. However, the absence of a standalone statistical software package to facilitate DDDM for large healthcare administrative databases has remained a significant gap. This paper introduces the R statistical software package CMHSU, available on the Comprehensive R Archive Network (CRAN), for the diagnosis of mental health (MH), substance use (SU), and their concurrent status (MHSU). The package implements DDDM using hospital and medical service physician visit counts along with maximum time span parameters for MH, SU, and MHSU diagnoses. A working example using a simulated real-world dataset is presented to examine various analytical aspects, including three key dimensions of MHSU detection based on the DDDM framework, as well as temporal analysis to demonstrate the package’s application for healthcare policymakers. Additionally, the limitations of the CMHSU package and potential directions for its future extension are discussed.
Keywords: Data-Driven Diagnostic Method (DDDM); Mental Health; Substance Use; Concurrent Status; R; Statistical Software; Healthcare Databases
@@ 1.2 North American Healthcare Administrative Databases
Country Level Hospital Data Physician Data Canada Federal Discharge Abstract Database (DAD) National Ambulatory Care Reporting System (NACRS) Canadian Institute of Health Information(CIHI) Physician Claims Data Canadian Management Information Database (CMDB) Province MED-ÉCHO (Quebec) data Ontario Health data Alberta Health data Medical Services Plan (MSP) Data (BC) Alberta Health Physician Claims Régie de l’assurance maladie du Québec (RAMQ) Ontario Health Insurance Plan (OHIP) Territory Yukon Hospital Corporation Dis-charge Data Northwest Territories Health Authority Records Yukon Health Care Insurance Plan Claims Northwest Territories Medical Billing Data United States Federal National Inpatient Sample (NIS) Medicare Provider Analysis and Review (MEDPAR) Veterans Affairs (VA) Inpatient Data Medicare Physician/Supplier Part B Claims Data National Ambulatory Medical Care Survey (NAMCS) Veterans Affairs (VA) Outpatient Data State State Inpatient Databases (SID) California Office of Statewide Health Planning and Development (OSHPD) New York Statewide Planning and Research Cooperative System (SPARCS) State Medicaid Claims Data All-Payer Claims Databases (APCDs: e.g., Massachusetts, Colorado) Territory Guam Memorial Hospital Authority Records American Samoa Dept. of Health Inpatient Data Guam Medicaid and CHIP Claims Data American Samoa Medicaid Claims1
@@ 1.3 Motivation
The DDDM has several advantages over other methods for the diagnosis of MHSU:
@@ 4.2 Strengths, Limitations and Future Work
    Empirical Validation: The simulation study in this paper uses self-generated simulated data. The optimal scenario would involve using a real-life empirical healthcare administrative database, which would significantly enhance the validity and usefulness of the results. However, key obstacles such as (1) Data Privacy and Security Regulations and (2) Access Restrictions and Bureaucratic Hurdles made it impossible to conduct the current analysis on such a database.
@@
    Geographical Adaptability: CMHSU is designed for detecting mental health status (MH), substance use status (SU), and their concurrent status (MHSU) within North American healthcare databases, based on the original DDDM developed by researchers in this region. However, it remains unclear whether these methods can be effectively adapted for use in non-North American healthcare systems, given potential differences in healthcare infrastructure, coding practices, and administrative data structures.
@@
    Temporal Analysis: The simulation study in this paper serves as a hypothetical demonstration of the package’s application. Its results can only be truly meaningful and applicable to real-world policymaking if based on a universally agreed-upon set of predefined parameters related to time span and visit definitions. However, no such universally accepted standard currently exists.
@@ 5 Conclusions
Acknowledgments. This paper is dedicated to the families and loved ones of those affected by the illicit drugs crisis in Canada and the United States. May this work serve as another steppingstone towards meaningful solutions that prevent further loss and offer healing to communities nationwide.

### v4

@@ Abstract
The concept of concurrent mental health and substance use (MHSU) and its detection in patients has garnered growing interest among psychiatrists and healthcare policymakers over the past four decades. Researchers have proposed various diagnostic methods, including the Data-Driven Diagnostic Method (DDDM), for the identification of MHSU. However, the absence of a standalone statistical software package to facilitate DDDM for large healthcare administrative databases has remained a significant gap. This paper introduces the R statistical software package CMHSU, available on the Comprehensive R Archive Network (CRAN), for the diagnosis of mental health (MH), substance use (SU), and their concurrent status (MHSU). The package implements DDDM using hospital and medical service physician visit counts along with maximum time span parameters for MH, SU, and MHSU diagnoses. A simulated real-world dataset incorporating fentanyl is presented to examine various analytical aspects, including three key dimensions of MHSU detection based on the DDDM framework, as well as temporal analysis to demonstrate the package’s application for healthcare policymakers. Additionally, the limitations of the CMHSU package and potential directions for its future extension are discussed.
Keywords: Data-Driven Diagnostic Method (DDDM); Mental Health; Substance Use; fentanyl; Concurrent Status; R; Statistical Software; Healthcare Databases
@@ 1.2 North American Healthcare Administrative Databases
Country Level Hospital Data Physician Data Canada ([Uncaptioned image]) Federal Discharge Abstract Database (DAD) National Ambulatory Care Reporting System (NACRS) Canadian Institute of Health Information(CIHI) Physician Claims Data Canadian Management Information Database (CMDB) Province MED-ÉCHO (Quebec) data Ontario Health data Alberta Health data Medical Services Plan (MSP) Data (BC) Alberta Health Physician Claims Régie de l’assurance maladie du Québec (RAMQ) Ontario Health Insurance Plan (OHIP) Territory Yukon Hospital Corporation Dis-charge Data Northwest Territories Health Authority Records Yukon Health Care Insurance Plan Claims Northwest Territories Medical Billing Data United States ([Uncaptioned image]) Federal National Inpatient Sample (NIS) Medicare Provider Analysis and Review (MEDPAR) Veterans Affairs (VA) Inpatient Data Medicare Physician/Supplier Part B Claims Data National Ambulatory Medical Care Survey (NAMCS) Veterans Affairs (VA) Outpatient Data State State Inpatient Databases (SID) California Office of Statewide Health Planning and Development (OSHPD) New York Statewide Planning and Research Cooperative System (SPARCS) State Medicaid Claims Data All-Payer Claims Databases (APCDs: e.g., Massachusetts, Colorado) Territory Guam Memorial Hospital Authority Records American Samoa Dept. of Health Inpatient Data Guam Medicaid and CHIP Claims Data American Samoa Medicaid Claims1
@@ 1.3 Motivation
The DDDM has several advantages over other methods for the diagnosis of MHSU as follows:
@@ 3.5 Temporal Analysis
Remark 3.2.
Given the nature of the healthcare administrative database and the study objectives, principal investigators have the flexibility to consider various temporal components for [Unit, Span] in the temporal analysis. In this study, our temporal analysis was conducted to examine monthly variations over a year, represented as [Month, Year]. Other potential examples include [Day, Month], [Week, Year], [Quarter, Year], and [Year, Decade].
@@ 4.2 Strengths, Limitations and Future Work
    Empirical Validation: The simulation study presented in this paper utilizes self-generated simulated data. Ideally, using a real-life empirical healthcare administrative database would significantly enhance the validity and applicability of the findings for healthcare policy implementations. However, key obstacles—including (1) data privacy and security regulations, (2) access restrictions and bureaucratic hurdles, and (3) selection bias—prevented the use of such database in the current analysis.
@@
    Geographical Adaptability: CMHSU is designed for detecting mental health (MH) status , substance use (SU) status , and their concurrent (MHSU) status within the North American healthcare databases, based on the fact that the original DDDM was developed by researchers in this region. However, it remains unclear whether these methods can be effectively adopted for use in the non-North American healthcare systems, given potential differences in healthcare infrastructure, coding practices, and administrative data structures. Despite all these issues, given availability of the essential data fields, a preliminary analysis is still possible (See Section 4.3).
@@
    Temporal Analysis: The simulation study in this paper serves as a hypothetical demonstration of the package’s application. Its results can only be truly meaningful and applicable to real-world policymaking if the DDDM parameters were based on a universally agreed-upon set of predefined values related to time span and visit definitions. However, no such universally accepted standard parameters currently exists among psychiatrists. Addressing this issue requires further discussions, research and a final systematic review and meta-analysis among various studies.
4.3 Summary of CMHSU Data Analysis Workflow
We conclude this section with a summary of the data analysis workflow, designed to facilitate a smooth user experience when applying the CMHSU package in their projects. The workflow comprises the following steps:
  1. (i)
    Compile CMHSU Data:
    Gather essential data fields, including $ClientID,VisitDate,Diagnostic_{H},and,\ Diagnostic_{P}.$
  2. (ii)
    Assess Scalability:
    • (a)
      Size: Execute data size scalability by applying the function splitfunction_id().
    • (b)
      Time: Execute temporal scalability by applying the function splitfunction_time().
  3. (iii)
    Define Analysis Parameters:
    Specify required parameters, such as n_MHH, n_MHP, and others as necessary.
  4. (iv)
    Examine Data Time Span:
    • (a)
      For data spans less than or equal to t_MHSU, apply the function MHSU_status_basic().
    • (b)
      For data spans exceeding t_MHSU, apply the function MHSU_status_broad().
  5. (v)
    Report Results:
    • (a)
      Frequency: Extract count data using the script SummarySampleMHSU_1.
    • (b)
      Proportion: Extract proportion data using the script SummarySampleMHSU_1.
  6. (vi)
    Temporal Interpretation:
    Interpret temporal patterns by applying the [Unit, Span] methodology illustrated in Section 3.5.
@@ 5 Conclusions
Acknowledgments. This paper is dedicated to the families and loved ones of those affected by the fentanyl and other illicit drugs crisis in Canada and the United States. May this work serve as another steppingstone towards meaningful solutions that prevent further loss and offer healing to communities in both nations ([Uncaptioned image]).